OpenStudy (anonymous):

Solve the equation: z^3=z+z* I'm super lost; been trying to expand using z=a+bi but without going anywhere. I also tried utilizing z^n=r^n(cos(alpha)+i*sin(alpha)) without not really knowing what to do. Please halp! :c

OpenStudy (anonymous):

Lol im studying tat right now

OpenStudy (anonymous):

Ill help you in this when I get a chance tomorrow, tag me in this so I can find it tomorrow

OpenStudy (anonymous):

Just got this one question; I know the solutions. Got no clue how to get there though.

OpenStudy (anonymous):

Does z* imply the conjugate?

OpenStudy (anonymous):

Indeedio.

OpenStudy (anonymous):

I was not sure how to write it out.

OpenStudy (anonymous):

okay, i will try.

OpenStudy (anonymous):

Thanks! :) Trying to come up with new stuff as we speak as well.

OpenStudy (anonymous):

Try, Z= a+bi. you get \[(a+ib)^3 = 2a\] Now expand. The terms with 'i' add up to zero. The rest equal to 2a.

OpenStudy (anonymous):

How do they add up to zero?

OpenStudy (anonymous):

I get a^3-2ab^2-ab^2-2a+(3a^2b-b^3)*i=0

OpenStudy (anonymous):

Because the imaginary part on both the sides should be equal. I will give the working in next post.

OpenStudy (anonymous):

Oh, right!

OpenStudy (anonymous):

\[a^3+3a^2bi-3ab^2-ib^3 =a+ib+a-ib \] \[(a^3-3ab^2)+i(3a^2b-b^3)=2a\] Now equating real and imaginary parts. \[a^3-3ab^2=2a \] \[3a^2-b^3=0\] Can you continue from here?

OpenStudy (anonymous):

This is precisely the step I'm at.

OpenStudy (anonymous):

a=(a^3-3ab^2)/2 and insert into (2) in the set. and pray to god it gets better(?).

OpenStudy (anonymous):

The last equation has b multiplied to a^2. sorry. I feel this is better. On simplifying second one,\[3a^2=b^2\] Now put b^2 in the first. You have a third degree equation in 'a'.

OpenStudy (anonymous):

I'll manage from here, thanks a lot! :)

OpenStudy (anonymous):

(I think!)

OpenStudy (anonymous):

I get the fator a(4a^2+1)=0

OpenStudy (anonymous):

Good, I got stuck up right there.

OpenStudy (anonymous):

But solving (4a^2+1)=0 yields wrong answer for a. :/

OpenStudy (anonymous):

a=+-i/2

OpenStudy (anonymous):

Looks like there was nothing wrong in the procedure. But the result is wrong.

OpenStudy (anonymous):

Indeed, maybe the method is not viable?

OpenStudy (anonymous):

If we only had a(a^2-2)=0 for some reason. :x

OpenStudy (anonymous):

ooooh.

OpenStudy (anonymous):

Aren't you dividing with zero at 3a^2b=b^3 --> 3a^2=b^2

OpenStudy (anonymous):

b may be zero.

OpenStudy (anonymous):

Yeah. :/

OpenStudy (anonymous):

If we solve using trigonometry, (try) \[\cos 3\theta =2 \cos \theta \] \[\sin 3\theta = 0\]

OpenStudy (anonymous):

Where did you go from?

OpenStudy (anonymous):

\[z= \cos \theta + i \sin \theta \]

OpenStudy (anonymous):

How cos(3Theta)? I'm stupidly bad at trigonometry. :c

OpenStudy (anonymous):

From de' moivre's theorem.

OpenStudy (anonymous):

Isn't z^n=r^n*(cos(Theta)+i*sin(Theta))?

OpenStudy (anonymous):

No, \[(\cos \theta + i \sin \theta)^n = \cos n \theta +i \sin n \theta\]

OpenStudy (anonymous):

Ok yeah, I follow to cos(3Theta)=2cos(Theta) now.

OpenStudy (anonymous):

Okay.do we get a answer this way?

OpenStudy (anonymous):

I got noooooo idea dude! :D Not sure where to go from cos(3Theta)=2cos(Theta) and sin(3*theta)=0

OpenStudy (anonymous):

arcsin(0)=0 --> Theta=0 (?)

OpenStudy (anonymous):

Which one is the theta?

OpenStudy (anonymous):

I'm super confused and got no idea what's happening atm.

OpenStudy (anonymous):

What are your thoughts? Have you even solved it perhaps?

OpenStudy (anonymous):

I am again getting a absurd result that cos is grater than 1. Before getting it I divided the equation by cos.

OpenStudy (anonymous):

:C

OpenStudy (anonymous):

I am not able to solve this. Not getting anything. Is there any solution for this?

OpenStudy (anonymous):

The solutions are z=0,z=-sqrt(2),z=sqrt(2)

OpenStudy (anonymous):

z=0 is obvious. I'll try. If you get a way, inform me.

OpenStudy (anonymous):

You know, the set of equations pan out correctly, perhaps they're solvable somehow?

OpenStudy (anonymous):

a^3-3ab^2=2a,3a^2b=b^3 that is

OpenStudy (anonymous):

Or perhaps that way is only solvable numerically.

OpenStudy (anonymous):

Super conundrum. :/

OpenStudy (anonymous):

Yeah, Using guess method we get the answers this way. \[b(3a^2-b^2) = 0\] \[a(a^2-3b^2-2)=0\] Let b=0 be the solution of first equation. Put it in the second assuming a is not zero. You get a = +-sqrt(2).

OpenStudy (anonymous):

I don't deal in guesstimating. :D

OpenStudy (anonymous):

The other possibilities are not possible, since we just now tried to do.(above)

OpenStudy (anonymous):

This is very solvable, I just don't know how. :(

OpenStudy (anonymous):

Try any other way, you will end up with a wrong answer. This is the only possible way. I couldn't see it until now. :(

OpenStudy (anonymous):

But you can't just assume a is not zero.

OpenStudy (anonymous):

But you get a is sqrt(2). So, it confirms that your assumption is right. You can then verify using the question.

OpenStudy (anonymous):

Of course it might be right, that doesn't make it right. As I said, this is 100% solvable without assumptions or numerical methods.

OpenStudy (anonymous):

Man I feel dumb. :s

OpenStudy (anonymous):

Okay, When you get the method, post it. As far as I know, we use either z=a+ib or the trigonometry. Keep trying..

OpenStudy (shubhamsrg):

I didn;t read all of the above, did you get get the answer ?

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

I know the answer, I've known it before I even asked here.

OpenStudy (anonymous):

It's the way to the answer I need. :/

OpenStudy (shubhamsrg):

what is the answer ?

OpenStudy (anonymous):

z=0,z=sqrt(2),z=-sqrt(2)

OpenStudy (shubhamsrg):

this means 3a^2 = b^2 should be wrong.

OpenStudy (shubhamsrg):

its true only for z=0. hmm

OpenStudy (anonymous):

Indeed it is(?).

OpenStudy (anonymous):

Arriving at 3a^2=b^2 you divide with b which is potentially 0.

OpenStudy (shubhamsrg):

(a+ib)^3 = 2a a^3 - b^3 i + 3ib a^2 - 3ab^2 = 2a (a^3 - b^3 - 3ab^2 -2a) + i ( 3b a^2 - b^3) =0 so b(3a^2 - b^2) = 0 b=0 is 1 solution plugging that in (a^3 - b^3 - 3ab^2 -2a) = 0, you get the required value of a. Only questions remains why we ignored 3a^2 - b^2 =0 hmm

OpenStudy (anonymous):

Isn't b=0 only part of a solution?

OpenStudy (shubhamsrg):

b=0 is a part of the solution.

OpenStudy (experimentx):

what is z* ? conjugate of z ?

OpenStudy (anonymous):

Yes! So sorry for the late reply, didn't think you would respond so quickly.

OpenStudy (experimentx):

\[ z^3 = 2 \Re[z] \] has to be something that if rotated by 3 times goes to real axis. \[ \large r^3 e^{i3\theta }= r^3 e^{in \pi }\]

OpenStudy (experimentx):

try solving this for other value than zero, might work but not sure

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I'm not sure what you mean or what direction I'm supposed to be heading in.

OpenStudy (anonymous):

0+n*2Pi=Theta gives r(r^2-2)=0 The solutions are supposed to be z=0,z=-sqrt(2),z=sqrt(2)

OpenStudy (anonymous):

The solutions to z^3=z+z* that is.""

OpenStudy (experimentx):

z=0 is trivial solution

OpenStudy (anonymous):

Sure

OpenStudy (experimentx):

\[ \theta = 0, 60^\circ, 120, ... \] http://www.wolframalpha.com/input/?i=%28r%29%5E3+%3D+2+r+cos%2860+degree%29 \[ z = \pm e^{i 2\pi/ 6}\]

OpenStudy (anonymous):

I'm sorry, I don't follow at all. :(

OpenStudy (experimentx):

\[ z = 0, \frac 1 2 + i \frac 1 2, -(\frac 1 2 + i \frac 1 2) \]

OpenStudy (experimentx):

woops!! where did it go wrong then

OpenStudy (anonymous):

This problem is killing me. (Thanks so much for trying to help me btw) Got no clue what mistake you did.

OpenStudy (experimentx):

it seems like \( \theta = 0 \) is the only solution http://www.wolframalpha.com/input/?i=%28E%5E%28I+2Pi%2F3%29%29%5E3+%3D+E%5E%28I+2Pi%2F3%29+%2B+E%5E%28-I+2Pi%2F3%29

OpenStudy (experimentx):

\[ r^3 = 2r \cos (0) = 2 r \] it gives you 0, +- sqrt(2)

OpenStudy (anonymous):

So that WAS it, god damn it problem. :D

OpenStudy (anonymous):

But how would you draw that conclusion algebraically?

OpenStudy (experimentx):

let \( z = r e^{i \theta }\) be your solution. since z^3 = 2 Re z = 2 r cos(theta), 3 theta = npi => theta = n pi/3 show that n=0 is only solution and solve for 'r'